40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 13 20:03:16 2018, johannes, Configuration, Computers, c1auxex replacement IMG_20180313_105154890.jpgIMG_20180313_133031002.jpg
    Reply  Wed Mar 14 23:58:30 2018, johannes, Configuration, Computers, c1auxex replacement 
       Reply  Mon Mar 19 14:39:09 2018, johannes, Configuration, Computers, c1auxex replacement 
Message ID: 13687     Entry time: Mon Mar 19 14:39:09 2018     In reply to: 13682
Author: johannes 
Type: Configuration 
Category: Computers 
Subject: c1auxex replacement 

[gautam, johannes]

The temperature control output channel for the XEND seismometer wasn't working properly. The EPICS channel existed, could be written to and read from, but no physical voltage was observed on the (confirmed properly) wired connector.

The Acromag DAC that outputs this channel was completely spare in the original scheme and does not serve any other channels at the moment. We found it to be unresponsive to ping from the host machine (reminder: the Acromags are on their own subnet with IPs 192.168.114.xxx connected to the secondary ethernet adapter of c1auxex), while all others returned the ping just fine. The modules have daisy-chained ethernet connections, and the one Acromag unit behind the unresponsive one in the chain was still responding to ping and its channels were working, so it couldn't have been a problem with the (ethernet) cabling.

Gautam and I power-cycled the chassis and server, which resolved the issue. The channel is now outputting the requested voltage on the Out1 BNC connector of the chassis (front). When I was setting up the whole system and did frequent rebooting and IP-redefinitions I have seen network issues arise between server and Acromags. In particular, when changing the network settings server-side, the Acromags needed to reboot occasionally. So this whole problem was probably due to the recent server-swap, as the chassis had not been power-cycled since.

 

During the debugging we also found that the c1psl2 channels were not working. This was because I had overlooked to update the epics environment variables for the modbus path defined in /cvs/cds/caltech/target/c1psl2/npro_config.cmd from the local installation /opt/epics/ (which doesn't exist on the new server anymore) to the network location /cvs/cds/rtapps/epics-3.14.12.2_long/. This has been fixed and the slow diagnostic PSL channels are recording again.

ELOG V3.1.3-